OpenStudy (anonymous):

Which of the following expressions is equal to (x+2)+(x-2)(2x+1)? a. 2x^-2x b. 2x^2-x c. 2x^2+x d. 4x^2+2x

OpenStudy (amoodarya):

(x+2)+(x-2)(2x+1)=(x+2)(1+(2x+1)) =(x+2)(2x+2) now multiply

OpenStudy (amoodarya):

=2(x+2)(x+1)